# Performance-lossless Black-box Model Watermarking

### v1

@@ 3.2.3. Defining Watermarking Protocol
Throughout this work, let $\lambda\in\mathbb{N}$ denote a security parameter, implicit input for all defined algorithms. A function $\epsilon$ is termed negligible if it approaches zero faster than any polynomial function. We represent a probabilistic polynomial-time algorithm as PPT. $API^{W}_{M}$ denotes the application programming interface of model $M$ after deploying watermarking module $W$, and when the existence of the inner watermarking module is unknown, it is represented as $API^{*}_{M}$. $W_{c,k}$ denotes the watermarking module that incorporates copyright information $c$ and key $k$ as internal parameters. The pair $(x,y)$ denotes a query-response pair for $API_{M}$, where $y=API_{M}(x)$, with $x\in X$ and $y\in Y$, where $X$ and $Y$ represent the input content space and response content space, respectively. $x_{*}$ represents a query $x$ that contains the message ***, and $y_{*}$ represents a response $y$ that contains the message ***, where *$*\in\{trigger,noise\}$, and $trigger,noise$ denote meaningful triggers and meaningless noise, respectively.
Definition 1. (Branch Backdoor) A branch backdoor is a backdoor planting algorithm that adds a backdoor path (i.e., malicious function) in parallel to the normal path within a program. This backdoor path can activate the malicious function under specified conditions. When this is used for the planting watermark function, it can also be referred to as a branch watermarking algorithm. Branch backdoor schemes adhere to an essential property:
- Inference process distortion-free: The entire backdoor is inserted as an external function on a branch independent of the original function. This means that, except for triggers, queries have invariance with respect to the original task.
@@
Definition 2. (Branch Backdoor-based Watermarking Protocol for Model) A branch backdoor-based watermarking protocol for model is a set of possibly probabilistic algorithms, BBWP = (KeyGen, Backdoor-Activate, Branch-Deploy, Backdoor-Collection, Verify).
@@
- Backdoor-Activate is the trigger generation algorithm, taking an input of arbitrary length query $x$ and key $k$ and producing a trigger version $x_{tr}$ for $x$. This corresponds to the trigger generator of the outer module.
- Branch-Deploy is the watermark deployment algorithm, taking the $API_{M}$, watermarking algorithm $W$, copyright information $c$, and key $k$ as inputs and generating the $API^{W_{c,k}}_{M}$ with encapsulated watermark functionality. The watermark-deployed $API^{BW}_{M}$ has both an original inference path $M$ and a backdoor watermark path $BW$, which is on a parallel branch.
- Backdoor-Collection is the copyright evidence collection algorithm, taking the $API^{*}_{S}$ of the suspicious model $S$ and trigger $x_{tr}$ as inputs and outputting the copyright evidence $(x_{tr},y_{*})$ for $API^{*}_{S}$, a key input for the Verify algorithm. This corresponds to the inner module.
- Verify is the copyright verification algorithm, taking copyright information $c$, key $k$, and query-response pair $(x_{tr},y_{*})$ concerning $API^{*}_{S}$ as inputs and outputting the whether the pair $(x_{tr},y_{*})$ is valid evidence. This corresponds to the evidence detector of the outer module. If $y_{*}$ is $y_{c}$, the evidence is valid; otherwise, it is invalid.
Definition 3. (API State) An API state represents the response status of an application programming interface when presented with different queries. The Service State indicates that the watermark function has not been activated, while the Forensic State signifies that it has been activated. APIs deployed with watermarking algorithms have two response states: Service State and Forensic State, whereas original APIs always remain in the Forensic State.
Definition 4. (Backdoor-Collection Algorithm) This is a supplementary definition of the Backdoor-Collection algorithm in BBWP. It consists of two PPT sub-algorithms, simply denoted as BW = (Detect, Prove):
@@
- Prove is the copyright evidence generation algorithm, taking the original response content $y_{noise}$ of $API_{M}$, copyright information $c$, and key $k$ as input and outputting the response content $y_{c}$ with embedded $c$.
@@
For $API^{W_{c,k}}_{M}$, the output of the Detect algorithm, $s(x_{noise})=0$, implies the non-activation of the watermark function executed by the Prove function. Thus, the API is in the Service State, and $s(x_{tr})=1$ implies activation of the Prove function, switching the API from the Service State to the Forensic State.
3.2.4. Performance-lossless
For no-box watermarking, the related concepts of “undetectabilit” and “distortion-free” were initially introduced in these works (Christ et al., 2023; Kuditipudi et al., 2023), respectively, representing the same underlying concept: watermark operations have a negligible impact on the quality of each output text. For black-box watermarks, similar concepts are typically referred to as “Functionality-preserving”, signifying that the watermark’s impact on the inference performance for queries other than triggers is negligible. These three definitions use different terminology but share a common concept, which we summarize as “performance-lossles” and provide the following definition.
Definition 6. (($\epsilon,\lambda$)-performance-lossless) We call a BBWP
@@ 3.2.4. Performance-lossless
($\epsilon,\lambda$)-performance-lossless if the deployed watermark function only has a negligible impact on the performance of the protected model. In other words, for all PPT algorithms $\mathcal{A}$, $\mathcal{A}$’s distinguishing advantage between $API^{W_{\cdot,KeyGen(1^{\lambda})}}_{M}$ and $API_{M}$ is negligible for the security parameter $\lambda$.
3.3. Provable Performance-lossless
@@
Inspired by undetectable backdoors (Goldwasser et al., 2022), we utilize secure cryptographic primitives to construct a provably performance-lossless watermarking module as the added branch, as shown in Figure 3. The performance-lossless property of branch watermarking is reduced to a message authentication code (MAC) algorithm, which meets Existential Unforgeability under Chosen Message Attack (EUF-CMA). To better understand the relationship between the unforgeability of MAC and the performance-lossless property of the BBWP, we review some cryptographic primitives and propose the concept of Unlearnable Trigger (UT) based on them. Finally, we prove the proposed is performance-lossless by reduction.
3.3.1. Cryptographic Primitive
First of all, let us review some cryptography knowledge.
@@ 3.3.1. Cryptographic Primitive
Definition 5. (PseudoRandom Function, PRF) For the function set $\mathcal{F}$, there is $\forall f\in\mathcal{F},f:\{0,1\}^{\lambda}\to\{0,1\}^{\lambda}$, and the output distribution is $\mathcal{P}_{\lambda}$. For any probabilistic polynomial time (PPT) discriminator $\mathcal{A}$, if $\mathcal{A}$ adaptively submits multiple inputs $x_{i}$ to $\mathcal{F}$ and obtains the output $y_{i}$, the advantages of distinguishing situations I and II based on the output $y_{i}$ is negligible, then $\mathcal{F}$ is said to be PRF on $\mathcal{P}_{\lambda}$:
$\textbf{Adv}^{PRF}_{\mathcal{A}}=\bigg{|}\mathop{Pr}_{\text{I: }y_{i}=f(x_{i}),f\leftarrow\mathcal{F}}\left[\mathcal{A}(\text{I})=1\right]-\mathop{Pr}_{\text{II: }y_{i}\leftarrow\{0,1\}^{\lambda}}\left[\mathcal{A}(\text{II})=1\right]\bigg{|}.$
@@
[…]
If the MAC algorithm satisfies the PRF security in Definition 1, it satisfies EUF-CMA, which is defined as follows.
[…]
$\mathop{Max}_{A\in\mathcal{A}(t,q)}\{\textbf{Adv}^{\textit{EUF-CMA}}_{\mathcal{A}}\}\leq\epsilon(\lambda).$
@@
3.3.2. Unlearnable Trigger
As shown in Figure 3, we construct the Trigger Generator and Trigger Detector sub-module in BBWP based on the MAC algorithm. If the MAC algorithm satisfies $(\epsilon,t,q,\lambda)$ security in Definition 7, the BBWP satisfies the Unlearnable Trigger (UT) property, which is defined as follows.
@@ 3.3.2. Unlearnable Trigger
$Pr\left[\text{ideal}\right]=Pr\left[Detect_{k}(x_{tr}^{\prime})=1\bigg{|}\begin{split}KeyGen(1^{\lambda})=k;\\ \mathcal{A}(\mathcal{O}^{API})=x^{\prime},x_{tr}^{\prime}\end{split}\right],$
$\textbf{Adv}^{UT}_{\mathcal{A}}=\bigg{|}Pr\left[\text{real}\right]-Pr\left[\text{ideal}\right]\bigg{|},$
[…]
@@
3.3.3. Proof
By definition in section 3.2.3 and 3.2.4, a BBWP seems to naturally satisfy performance-lossless because the backdoor is introduced into the parallel branch as an external function independent of the model inference process, which means that the model will maintain its initial inference results unless the branch is activated by a trigger.
However, practical implementation needs some care. The Trigger Detector of the BBWP should have a negligible false positive rate (i.e., soundness) and a negligible false negative rate (i.e., completeness). The significance of an unlearnable trigger distribution is that even if an attacker deliberately constructs a query, they cannot bypass the trigger detector. Therefore, if the BBWP satisfies the $(\epsilon,t,q,\lambda)$-UT property in Definition 7, the Trigger Detector sub-module satisfies completeness and soundness. Then, a BBWP based on a EUF-CMA, i.e., $(\epsilon,t,q,\lambda)$ secure, MAC scheme is a $(\epsilon,\lambda)$-performance-lossless BBWP, and the proof is as follows:
@@ 3.3.3. Proof
Completeness. This property is satisfied based on a proof of correctness of the MAC scheme. Tags generated by a Mac function under the same symmetric key and the same message can always be the same. Therefore, as shown in Figure 3, triggers generated by the Trigger Generator can always be detected by the Trigger Detector under the symmetric key without considering the possibility of an attacker distorting the triggers.
[…]
Figure 3. Construction satisfying the unleanable trigger property.

### v2

@@ III-C2 Defining Watermarking Protocol
Throughout this work, let $\lambda\in\mathbb{N}$ denote a security parameter, which is an implicit input for all the defined algorithms. A function $\epsilon$ is termed negligible if it approaches zero faster than any polynomial function. We represent a probabilistic polynomial-time algorithm as PPT. $API^{W}_{M}$ denotes the API of model $M$ after deploying watermarking module $W$, and when the existence of the inner watermarking module is unknown, it is represented as $API^{*}_{M}$. $W_{c,k}$ denotes the watermarking module that incorporates copyright information $c$ and key $k$ as internal parameters. The pair $(x,y)$ denotes a query-response pair for $API_{M}$, where $y=API_{M}(x)$, with $x\in X$ and $y\in Y$, where $X$ and $Y$ represent the input content space and response content space, respectively. $x_{*}$ represents a query $x$ that contains the message *, and $y_{*}$ represents a response $y$ that contains the message *, where $*\in\{tri,noi\}$, and $tri,noi$ denote meaningful triggers and meaningless noise, respectively.
@@
Definition 1. (Branch Backdoor-based Watermarking Protocol for Model) A branch backdoor-based watermarking protocol for model is a set of possibly probabilistic algorithms, BranchWM = (KeyGen, TriggerGen, Branch, Deloy, Verify).
@@
  • TriggerGen is the trigger generation algorithm, taking an input of arbitrary length query $x$ and key $k$ and producing a trigger version $x_{tri}$ for $x$. This corresponds to the trigger generator of the outer module.
  • Branch is the watermarking algorithm, taking the $API^{*}_{S}$ of the suspicious model $S$ and trigger $x_{tri}$ as inputs and outputting the copyright evidence $(x_{tri},y_{*})$ for $API^{*}_{S}$, which is a key input for the Verify algorithm. This corresponds to the inner module.
  • Deploy is the watermarking deployment algorithm, taking the $API_{M}$, watermarking algorithm $W$, copyright information $c$, and key $k$ as inputs and generating the $API^{W_{c,k}}_{M}$ with encapsulated watermarking functionality. The watermarking-deployed $API^{BW}_{M}$ has both an original inference path $M$ and a backdoor watermarking path $BW$, which is as a parallel branch.
  • Verify is the copyright verification algorithm, taking copyright information $c$, key $k$, and query-response pair $(x_{tri},y_{*})$ concerning $API^{*}_{S}$ as inputs and outputting whether the pair $(x_{tri},y_{*})$ is valid evidence. This corresponds to the evidence detector of the outer module. If $y_{*}$ is $y_{c}$, the evidence is valid; otherwise, it is invalid.
Definition 2. (API State) An API state represents the response status of an API when presented with different queries. The Service State indicates that the watermarking function has not been activated, while the Forensic State signifies that it has been activated. APIs deployed with watermarking algorithms have two response states—Service State and Forensic State—whereas original APIs always remain in the Service State.
Definition 3. (Branch Algorithm) This is a supplementary definition of the Branch algorithm in BranchWM. It consists of two PPT sub-algorithms, simply denoted as BW = (Detect, Prove):
@@
  • Prove is the copyright evidence generation algorithm, taking the original response content $y_{noi}$ of $API_{M}$, copyright information $c$, and key $k$ as input and outputting the response content $y_{c}$ with embedded $c$.
@@
For $API^{W_{c,k}}_{M}$, the output of the Detect algorithm, $s(x_{noi})=0$, implies the non-activation of the watermarking function executed by the Prove function. Thus, the API is in the Service State, and $s(x_{tri})=1$ implies activation of the Prove function, switching the API from the Service State to the Forensic State.
Definition 4. (($\epsilon,\lambda$)-correct) We call a BranchWM ($\epsilon,\lambda$)-correct if a BranchWM satisfy the following formulas:
$\textbf{Verify}(k,(\textbf{API}^{W}_{M}(k,x_{tri}),x_{tri}))=1,$ (1)
$\textbf{Verify}(k,(\textbf{API}_{M}(k,x_{tri}),x_{tri}))=0,$ (2)
$\textbf{Verify}(k,(\textbf{API}^{W}_{M}(\neg k,x_{tri}),x_{tri}))=0,$ (3)
where $k$ is the output of $KeyGen(1^{\lambda})$ and $\neg k$ is a random sequence whose length is equal to the length of $k$. $x$ is a normal sample and $x_{tri}$ is the output of $TriggerGen(x,k)$.
Figure 2: There are two modules in the IP verification protocol. The outer module is deployed outside the model API, responsible for generating triggers to obtain model copyright evidence and verifying the evidence. The inner module is added as a branch in the model API, responsible for checking whether the input is a trigger and generating the copyright evidence.
III-C3 Definition of Performance-lossless
We give the definition of performance-lossless for model watermarking, which intuitively means that the watermarking task does not affect the original task of the model.
@@ III-C3 Definition of Performance-lossless
Definition 5. (($\epsilon,t,q,\lambda$)-performance-lossless) We call a BranchWM ($\epsilon,t,q,\lambda$)-performance-lossless if the deployed watermarking function only has a negligible impact on the performance of the protected model. In other words, for any PPT algorithm $\mathcal{A}$, after $q$ queries on $API^{W_{\cdot,KeyGen(1^{\lambda})}}_{M}$ and TriggerGen Oracle within time $t$, $\mathcal{A}$’s distinguishing advantage between $API^{W_{\cdot,KeyGen(1^{\lambda})}}_{M}$ and $API_{M}$ is negligible for the security parameter $\lambda$.
Since this paper does not consider adversaries with white-box capabilities, the performance-lossless property defaults to black-box performance-lossless.
III-D Provable Performance-lossless Watermarking from Secure Cryptographic Primitive
@@
Inspired by undetectable backdoors [36], we utilize an MAC algorithm, which meets Existential Unforgeability under Chosen Message Attack (EUF-CMA), to construct the trigger generator and detector sub-modules. As shown in Figure 3, the authentication code of the original sample is used as the triggering pattern, whose detection efficiency is better than that of the digital signature scheme based on asymmetric encryption technology. To better understand the relationship between the unforgeability of the MAC scheme and the performance-lossless property of the BranchWM, we review some cryptographic primitives and propose the concept of Unaccessible Trigger (UT) based on them, whose pattern meets Distributional Unaccessibility under Chosen Sample Attack (DUL-CSA). Then, we use the MAC scheme to construct the forensic trigger which is named MAC-based unaccessible and fast trigger, i.e., MUFT. Finally, we prove that the proposed BranchWM is performance-lossless by reduction.
III-D1 Cryptographic Primitive
First, let us review some cryptography knowledge.
@@ III-D1 Cryptographic Primitive
Definition 6. (PseudoRandom Function, PRF) For the function set $\mathcal{F}$, there is $\forall f\in\mathcal{F},f:\{0,1\}^{\lambda}\to\{0,1\}^{\lambda}$, and the output distribution is $\mathcal{P}_{\lambda}$. For any probabilistic polynomial time (PPT) discriminator $\mathcal{A}$, if $\mathcal{A}$ adaptively submits multiple inputs $x_{i}$ to $\mathcal{F}$ and obtains the output $y_{i}$, the advantage of distinguishing situations I and II based on the output $y_{i}$ is negligible, then $\mathcal{F}$ can be said to be a PRF on $\mathcal{P}_{\lambda}$:
@@
$\textbf{Adv}^{PRF}_{\mathcal{A}}=\bigg{|}\mathop{Pr}\left[\mathcal{A}(\text{I})=1\right]-\mathop{Pr}\left[\mathcal{A}(\text{II})=1\right]\bigg{|}.$ (6)
  • KeyGen is the key generation algorithm, uniformly outputting a key $k$ of length $\lambda$ according to the selected security parameter $\lambda\leftarrow\textit{Key}(1^{\lambda})$;
  • Mac is the message encoding algorithm, taking $k\in\{0,1\}^{\lambda}$ and $M\in\{0,1\}^{L}$ as inputs and outputting $\textit{tag}\in\{0,1\}^{l}$, where tag is the message authentication code corresponding to $M$, that is, $\textit{tag}\leftarrow\textit{Mac}_{k}(M)$;
  • Veri is the message verification algorithm, taking $k\in\{0,1\}^{\lambda}$, $\textit{tag}\in\{0,1\}^{l}$, and $M\in\{0,1\}^{L}$ as inputs and outputting one bit $b\in\{0,1\}$; $\forall k$ and $M$, there is $\textit{Veri}_{k}(\textit{Mac}_{k}(M),M))=1$.
If the MAC scheme satisfies the PRF security in Definition 6, it satisfies EUF-CMA, which is defined as follows.
Definition 8. (Existential Unforgeability under Chosen Message Attack, EUF-CMA) For an MAC algorithm MAC = (Key, Mac, Veri), any PPT adversary $\mathcal{A}$ can select message $M_{i}\in\{0,1\}^{l}$ and query the MAC for $\textit{tag}_{i}=\textit{Mac}_{k}(M_{i})$ within time $t$. After $q$ queries, the MAC is said to satisfy EUF-CMA security, if the advantage that $\mathcal{A}$ outputs $(M^{\prime},\textit{tag}^{\prime})$ where $M^{\prime}\notin\{M_{1},\cdots,M_{q}\}$, $\textit{Veri}_{k}(\textit{tag}^{\prime},M^{\prime})=1$, is negligible:
$\mathop{Max}_{A\in\mathcal{A}(t,q)}\{\textbf{Adv}^{\textit{EUF-CMA}}_{\mathcal{A}}\}\leq\epsilon(\lambda).$ (7)
@@
Figure 3: We use the MAC scheme which meets EUF-CMA to construct the Trigger Generator and Trigger Detector sub-modules. Such a construction generates the MUFT that satisfies the DUL-CSA property.
III-D2 Unaccessible Trigger
We construct the Trigger Generator and Trigger Detector sub-modules in the BranchWM based on the MAC scheme. If the MAC scheme satisfies $(\epsilon,t,q,\lambda)$ security in Definition 8, the trigger pattern of BranchWM satisfies the Distributional Unaccessibility under Chosen Sample Attack (DUL-CSA), which is defined as follows.
Definition 9. (Distributional Unaccessibility under Chosen Sample Attack, DUL-CSA) For a BranchWM = (KeyGen, TriggerGen, Branch, Deploy, Verify), any PPT adversary $\mathcal{A}$ in the “real” world can select sample $x_{i}\in\{0,1\}^{l}$ and query $\mathcal{O}^{TG}$ (i.e., the trigger generator of the protocol’s outer module) for $x_{i,trigger}\leftarrow\textit{TriggerGen}_{k}(x_{i})$ within time $t$. After $q$ queries, BranchWM is said to satisfy $(\epsilon,t,q,\lambda)$-DUL-CSA property, if the advantage, that $\mathcal{A}$ output $(x^{\prime},x_{tri}^{\prime})$ where $x^{\prime}\notin\{x_{1},\cdots,x_{q}\}$, $\textit{Detect}_{k}(x_{tri}^{\prime},)=1$, is negligible:
@@ III-D2 Unaccessible Trigger
$Pr(\text{real})=Pr(Detect_{k}(x_{tri}^{\prime})=1|\text{I}),$ (10)
$Pr(\text{ideal})=Pr(Detect_{k}(x_{tri}^{\prime})=1|\text{II}),$ (11)
$\textbf{Adv}^{UT}_{\mathcal{A}}=\bigg{|}Pr(\text{real})-Pr(\text{ideal})\bigg{|},$ (12)
$\mathop{Max}_{A\in\mathcal{A}(t,q)}\{\textbf{Adv}^{\textit{UT}}_{\mathcal{A}}\}\leq\epsilon(\lambda).$ (13)
@@
Then, the trigger pattern of the B ran ch WM can be called a $(\epsilon,t,q,\lambda)$-secure trigger pattern. We refer to such triggers as unaccessible triggers. Because they are constructed from the MAC scheme, we also call them MAC-based unaccessible and fast triggers, i.e., MUFTs, where the “fast” is discussed in Section III-E3.
III-D3 Proof
The idea of multi-branch structure separates the model’s original task and forensic task. Unless a branch is activated by a trigger, the model will maintain its initial inference results. This means that the ideal model fidelity is determined by the reliability of the trigger mechanism; that is, the trigger detector of BranchWM should have a negligible false positive rate (i.e., robustness) and a negligible false negative rate (i.e., completeness). The significance of an unaccessible trigger pattern is that even if an attacker deliberately constructs a query, they cannot activate the trigger detector without the correct key used in trigger generation. Therefore, if the trigger pattern of BranchWM satisfies ($\epsilon,t,q,\lambda$)-DUL-CSA in Definition 9, the trigger detector submodule satisfies completeness and soundness. Then, a BranchWM constructed from a MAC scheme that meets ($\epsilon,t,q,\lambda$)-EUF-CMA is a ($\epsilon,t,q,\lambda$)-performance-lossless BranchWM, which is proven as follows:
@@ III-D3 Proof
Completeness. This property is satisfied based on a proof of correctness of the MAC scheme. Tags generated by a Mac function under the same symmetric key and the same message can always be the same. Therefore, as shown in Fig 3, triggers generated by the Trigger Generator can always be detected by the Trigger Detector under the symmetric key without considering the possibility of an attacker distorting the triggers.
Soundness. This property is reduced to the security of the MAC scheme. If there exists a PPT algorithm $\mathcal{A}$ that can generate a trigger query $x^{\prime}_{{trigger}}$ without knowledge of the symmetric key $k$ and successfully obtain $\textit{Detect}_{k}(x_{tri}^{\prime},)=1$, then $\mathcal{A}$ can be used to forge a valid tag on $x^{\prime}$ about $k$. This contradicts the DUL-CSA security of the trigger pattern, thereby contradicting the EUF-CMA security of MAC. Hence, without the knowledge of $k$, the probability that trigger queries can be successfully forged is negligible. Not to mention the possibility of triggering a forensic state with unintentional queries.
III-E Comparison of Secure Cryptography Primitives
III-E1 Secure Hash Function
A hash function $h$ is an algorithm that maps arbitrary length data to fixed length data. [^1] This is an informal definition of an individual hash function, not a hash function family. The hash function used in practice can be regarded as a hash function with determined parameters in the hash function family. For any input data, a hash function can generate a unique hash value. Generally speaking, a secure hash function has the following properties:
  • Preimage resistance: For any given hash output $y$, it is computationally infeasible to find a corresponding message input $x$ such that $h(x)=y$.
  • Second preimage resistance: For any given input $x$, it is computationally infeasible to find an input $x^{\prime}$ that is different from x such that $h(x)=h(x^{\prime})$.
  • Collision resistance: It is computationally infeasible to find two different inputs $x$ and $x^{\prime}$ such that $h(x)=h(x^{\prime})$.
III-E2 EUF-CMA Signature
A digital signature is a cryptographic technology used to verify and ensure the origin and integrity of digital information. It is based on asymmetric encryption and usually includes three algorithms: key generation, signing, and verification. To ensure security, existential unforgeability is an important security property of a digital signature scheme:
  • Existential unforgeability: Without holding the private key, an adversary with polynomial computing power, after querying a certain number of (message, signature) pairs, is still unable to construct a valid signature for a message $m$ that has not been previously queried for any corresponding signature.
Table I: The runtime cost of the verification time between different cryptographic primitives. We chose one hash function (i.e., SHA512), one MAC scheme (i.e., HMAC_SHA512), and two digital signature schemes (i.e., ECDSA and FALCON) according to the recommendations of NIST [52, 53, 54, 55]. We use SHA512 as the baseline to record the ratios of HMAC, ECDSA and FALCON under the 1024-bit, 163-bit and 10248-bit key lengths, respectively. FALCON_512 is its basic security version.
Method Key Size (bit) Output Size (bit) Time Rate
SHA512 /// $512$ $1(0.0114ms)$
HMAC_SHA512 $1024$ $512$ $\bm{1.2020_{\pm 0.0784}}$
ECDSA_SHA512 $163$ $326$ $64.6781_{\pm 2.1548}$
FALCON_512 $10248$ $5520$ $5.2550_{\pm 0.1899}$
To improve security, the signing algorithm is usually a probabilistic algorithm that introduces randomness into the signature generation process so that different signatures can be obtained when the same message is signed multiple times. The above security is a kind of weak existential unforgeability [^2] The existence unforgeability of MAC defined in this paper is weak existence unforgeability., and the adversary’s assumption is that it has never observed any signature of the message $m$. Accordingly, strong existential unforgeability means that after querying the signature of the message $m$, the attacker still cannot generate a valid new signature for it.
III-E3 Hash vs MAC vs Signature
Runtime Cost. Goldwasser et al. [36] used a strong unforgeable digital signature scheme to construct the trigger mechanism of the backdoor and further combined it with the Cook-Levin theorem to achieve the non-replicability (i.e., unforgeability) of the trigger in a white-box model situation. However, as shown in Table I, digital signature schemes based on asymmetric cryptography will introduce heavier operating overhead in the service state of the model API than hash functions and MACs. [^3] FALCON is a lattice-based post-quantum signature algorithm. Although it achieves high enough security with a short signature length and fast speed in the cryptograhic field, this signature size is still unacceptable as a trigger pattern of the language model watermarking. In addition, in the case of the white-box model, the model will face an easier but more severe attack—the adversary can identify and remove the backdoor or watermark based on structural anomalies. The benefits of white-box unforgeability are not sufficient to extend the capability of an adversary from the black-box to the white-box.
Trigger Pattern Space. Strong unforgeable digital signatures can generate multiple trigger patterns for the same original sample $x$, making the trigger pattern space larger. However, a large pattern space is not necessary. A large sample space can guarantee a large trigger space. It is feasible to generate a unique trigger pattern for $x$ based on a hash function or MAC and record $x$ to achieve a one-time trigger.
Security. Hash functions are the most efficient of the three cryptographic primitives. If a secure hash function is used to construct the trigger mechanism of the branch backdoor-based model watermarking, the running overhead introduced by the deployed watermark is small when the model API is in the service state. However, non-key-driven hash functions make it hard to protect the model without keeping the hash function secret.